A partnership has the following capital balances:

Comprix (50% of gains and losses) $ 100,000
Heflin (25%) 200,000
Kaplan (25%) 300,000

Mahar is going to pay a total of $202,500 directly to these three partners to acquire a 25 percent ownership interest from each. Goodwill is to be recorded. What is Heflin’s capital balance after the transaction?

Multiple Choice

  • $150,000

  • $175,000

  • $200,000

  • $189,375
Answer #1

Present capital = 100,000+200,000+300,000 = 600,000

Implied capital = 202,500/25% = 810,000

Goodwill = Implied capital - Present capital

= 810,000 - 600,000 = 210,000

Goodwill allocated to Heflin = 210,000*25% = 52,500

Capital of Heflin after goodwill allocation = 200,000 + 52,500 = 252,500

Heflin capital after the transaction = 252,500 - 25% given up

= $189,375